Please help me solve this: find three consecutive even integers such that the sum of the first at the third is 12 less than four times the second
Answer by sudhanshu_kmr(1152)


Let even integers are x, x+2 and x+4.
x + (x+4) = 4(x+2) - 12
=> 2x + 4 = 4x + 8 -12
=> 2x + 4 = 4x -4
=> 2x -4x = -8
=> -2x = -8
=> x = 4
integers are 4, 6 and 8.